I need to know what the ? next number is in these two sequences:
2, 7, 26, 101, 400, ?
7, 21, 8, 72, 9, ?

Answer by Edwin McCravy(20056)

  2*4 =    8, then    8 - 1 =    7
  7*4 =   28, then   28 - 2 =   26
 26*4 =  104, then  104 - 3 =  101
101*4 =  404, then  404 - 4 =  400 
400*4 = 1600, then 1600 - 5 = 1595

7
7*3 = 21
8
8*3*3 = 72
9
9*3*3*3 = 243

Edwin